OpenStudy (anonymous):

A cookie jar contains 3 chocolate chip cookies, 4 oatmeal raisin cookies, and 5 sugar cookies.Event A = drawing a chocolate chip cookie on the first draw Event B = drawing an oatmeal raisin cookie on the second drawEr two cookies are drawn from the jar, one after the other and not replaced, what is P(A and B) expressed in simplest form?

OpenStudy (anonymous):

\[P(A)=\frac{3}{12}=\frac{1}{4}\] \[P(B|A)=\frac{4}{11}\] reasoning that there are 11 cookies left and of those 4 are oatmeal (since first one is raisin) then \[P(A\cap B)=P(A)\times P(B|A)=\frac{1}{4}\times \frac{4}{11}\]

OpenStudy (anonymous):

thanks